OpenStudy (watsup232):

A common inhabitant of human intestines is the bacterium Escherichia coli, named after the German pediatrician Theodor Escherich, who identified it in 1885. A cell of this bacterium in a nutrient-broth medium divides into two cells every 20 minutes. The initial population of a culture is 40 cells. Find the Growth Rate.

likeabossssssss (likeabossssssss):

\(\bbox[5pt,midnightblue]{\LARGE\color{yellow}{\star ^\bigstar Welcome~to~~os~ ^\bigstar \star}}\)

OpenStudy (dumbcow):

Population doubles every 20 min, so when t =20, population = 40*2 exponential function will look like this: \[\large 40 (2^{\frac{t}{20}})\] The growth rate comes from the following exponential form \[(1+r)^t\] where "1+r" is the base of the exponential function set equal the bases and solve for r \[\rightarrow 1+r = 2^{\frac{1}{20}}\] \[r = 2^{\frac{1}{20}} - 1\]